Lindon Company is the exclusive distributor for an automotive product that sells for $60 per unit and has a CM ratio of 35%. The company’s fixed expenses are $226,800 per year. The company plans to sell 14,500 units this year. Required:

1. What are the variable expenses per unit?

2. Use the equation method: a. What is the break-even point in unit sales and in dollar sales? b. What amount of unit sales and dollar sales is required to earn an annual profit of $63,000? c. Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $3 per unit. What is the company’s new break-even point in unit sales and in dollar sales? .

3. Repeat (2) above using the formula method. a. What is the break-even point in unit sales and in dollar sales? b. What amount of unit sales and dollar sales is required to earn an annual profit of $63,000? c. Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $3 per unit. What is the company’s new break-even point in unit sales and in dollar sales?
Answer #1

1) variable cost per unit = 60*65% = 39 per unit

2a) Profit = Sales-variable cost-Fixed cost

0 = 60X-39X-226800

21X = 226800

X(Break even unit) = 226800/21 = 10800 Units

Break even sales = 10800*60 = $648000

2b) Profit = Sales-Variable cost-Fixed cost

63000 = 60X-39X-226800

21X = 289800

X(Required unit) = 13800 Units

Required sales = 13800*60 = 828000

2c) Break even unit :

Profit = Sales-Variable cost-Fixed cost

0 = 60X-36X-226800

24X = 226800

X(Break even unit) = 9450 Units

Break even sales = 9450*60 = 567000

3a) Break even unit = Fixed cost/Contribution margin per unit = 226800/60-39 = 10800 Units

Break even sales = 10800*60 = 648000

Required unit = (226800+63000)/21 = 13800 Units

required unit sales = 13800*60 = 828000

New break even unit = 226800/24 = 9450 Units

New break even sales = 9450*60 = 567000